OpenStudy (anonymous):

a forester planted 3/13 acre in seedlings in the morning and 12/13 acre in the afternoon. if 4/13 acre of seedlings were destroyed by frost. how many acres remained?

OpenStudy (anonymous):

Total = 3/13+12/13-4/13=11/13 acre

OpenStudy (anonymous):

why thank you!:)